OpenStudy (anonymous):

need help please. I don't understand. Draw the graph of relation defined by y^2=3x -first, find the vertex of the equation -select 1 pt above the axis of symmetry and 1 pt below thanks !!

OpenStudy (anonymous):

it's a parabola with vertex at the origin, opening towards the positive x axis...

OpenStudy (anonymous):

i still don't understand. vertex is (0,0) . how do i get the other points ?

OpenStudy (anonymous):

pick any x coordinate.... say x=1/3. then y = ???

OpenStudy (anonymous):

for any positive x value you pick, you will have two y-values... for example, when x=1/3, then \(\large y^2=3(\frac{1}{3})\rightarrow y=\pm1 \) here are your two points that correspond with x=1/3: (1/3, 1) and (1/3, -1) keep doing that for several more x values and you'll see you have a parabola opening towards the right with the vertex at the origin.

OpenStudy (anonymous):

got it ! thanks !